### v1

@@ 3.3 Training
Inspired by recent advances (LI et al., 2024), we introduce the Progressive Paradigm Training (PPT) strategy, which enables LLMs to gradually master an increasing variety of reasoning paradigms. As shown in Figure 3 (b), this training framework enables the model to progressively expand its reasoning capability by introducing different types of reasoning data in stages. Each training stage uses a different combination of reasoning paradigms. In the stage ➀, given the dominance of NL in the pre-training data of language models (Dubey et al., 2024), we create Numina-CoT as an initialized teaching stage. This is done by modifying the original Numina-CoT dataset (LI et al., 2024) according to our universal text template, which enables the model to learn to use NL to solve complex mathematical problems. Based on the question $x$, the model performs reasoning $\tau_{\text{NLR}}$ and generates the answer $y$. The generated sequence is $z=[x]\tau_{\text{NLR}}y$, where $[x]$ represents the tokens of the inputs. For simplicity, we first consider the loss function for a single sample:
@@
where $\theta$ represents the model parameters, $z_{t}$ is the $t$-th token in the generated sequence, and $z_{<t}$ indicates all tokens before the $t$-th token in the generated sequence. In stage ➁, considering a certain proportion of code corpora in the pre-training data, we expand the training dataset to include two paradigms: NLR and AR. Similar to Numina-CoT, we modify the original Numina-TIR to create Numina-TIR. With this modified dataset, the generated sequence is $z=[x]\tau_{\text{NLR}}\tau_{\text{AR}}y$. After this stage, the model can handle problems that require precise answers. Finally, in the stage ➂, we further expand the training data to three reasoning paradigms by utilizing the MPM dataset, where $z$ is $[x]\tau_{\text{NLR}}\tau_{\text{AR}}\tau_{\text{SR}}y$. After full PPT stages, the trained CoR-Math-7B model not only masters NLR and AR but also can perform rigorous logical SR.
@@ 3.4 Inference
We introduce an inference method that combines variable reasoning depth and sequential multi-paradigm sampling to solve comprehensive mathematical tasks, enabling the model to adapt its reasoning process based on task-specific requirements.
Prompts with variable reasoning depth. In the zero-shot inference phase, the CoR-Math-7B exhibits proficiency in multi-paradigm reasoning. This proficiency allows for the implementation of multi-paradigm reasoning with adaptable reasoning depths, achieved by adjusting prompts according to the specific demands of a given task. Initially, the model is prompted to conduct NLR, thereby activating a wide range of knowledge patterns associated with natural language from the pre-training corpus. Subsequently, the inference method is tailored to the specific problem type. For instance, in the context of theorem proving, the model transitions to SR, which is better suited for formal deduction. Given that the output of the model during SR is structured in Lean 4, the relevant proof segment can be extracted as the final solution. In cases of arithmetic computation, the model first performs NLR, followed by SR to strengthen logical coherence, and ultimately employs AR for precise calculations. A summary box is utilized to present the final results. This adaptable paradigm effectively captures the distinct reasoning patterns inherent in different paradigms and demonstrates flexibility in accommodating a variety of scenarios.
Sequential Multi-Paradigm Sampling (SMPS). Instead of token-level sampling based on tree structures (Qiu et al., 2024; Xiong et al., 2024) within single-paradigm reasoning paths, we purpose sequential paradigm-level sampling method, named SMPS. This approach allows the model to generate outputs by sequentially sampling across different paradigms. For instance, in a two-paradigm reasoning scenario, the model first instantiates $J$ distinct paths for the initial reasoning paradigm $\tau_{1}$.
@@
Subsequently, for each of these $\tau_{1}$ paths, the model instantiates $K$ possible paths for the secondary reasoning paradigm $\tau_{2}$
@@
In summary, the SMPS method leverages the combinatorial expansion of reasoning paths to explore a diverse paradigm-based solution space comprehensively. This structured approach enhances the robustness and depth of the reasoning results.
@@ 4.1 Evaluation Setup
Datasets. To evaluate our model’s mathematical reasoning capabilities comprehensively, we utilize a suite of widely recognized and representative benchmark datasets. These datasets encompass a diverse range of mathematical reasoning tasks, including arithmetic computation and theorem proving. The arithmetic computation ability is evaluated on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) datasets. These datasets span a spectrum of difficulty levels, from elementary arithmetic to mathematical competitions. Theorem proving ability is evaluated on the MiniF2F test set (Zheng et al., 2022), which features mathematical problems of Olympiad difficulty.
Metrics. Accuracy serves as the primary metric for performance evaluation across all datasets. In detail, we follow the widely-used CoT settings (Wei et al., 2022) to obtain solutions for arithmetic problem tasks, with numerical answers rounded to the nearest integer. The SymPy library [^1] https://www.sympy.org/ is utilized for parsing and evaluating mathematical expressions. To address potential variations in numerical representation, the model is instructed to explicitly state the final answer in the conclusion, following the approach in Numina models (LI et al., 2024). For theorem proving datasets, we follow the recent advances (Xin et al., 2024) to adapt the publicly available MiniF2F benchmark for the Lean 3 theorem proving environment to the Lean 4 environment [^2] https://github.com/rahul13613/miniF2F-lean4. These adjustments ensure the use of the most up-to-date formal verification tools. The pass@$N$ metric evaluates the capacity of the model to generate correct proofs within $N$ attempts. Specifically, our CoR-Math-7B employs SMPS, which operates by performing NLR followed by SR. Therefore, the sample budget $N$ is $N_{\text{NLR}}\times N_{\text{SR}}$. Detailed settings of all models are provided in Section A.5.
@@ 4.2 Implementation Details
We fine-tuned widely-used DeepSeek-Math-Base 7B (Shao et al., 2024) and Llama-3.1 8B (Dubey et al., 2024) models, employing our PPT method on MPM dataset. Unless otherwise specified, CoR-Math-7B model is based on DeepSeek-Math-Base 7B. The details are available in Section A.3.
@@ 4.3 Baselines
This study examines three categories of baseline models, with results reported using CoT prompting.
@@
General-purpose mathematical models: Considering the general applicability of our CoR-Math-7B model, we present the SOTA general-purpose mathematical models, including Mustard (Huang et al., 2024), DeepSeek-Math (Shao et al., 2024), InternLM-Math (Ying et al., 2024b), Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Llemma (Azerbayev et al., 2024).
@@
Task-specific mathematical models: We consider several expert models on mathematical optimization for specific tasks. For expert models on arithmetic computation, we include several models that undergo specialized reinforcement in arithmetic computation to achieve enhanced performance, such as more data and tool invocation. These models encompass Qwen2.5-Math (Yang et al., 2024), WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2024), DART-Math (Tong et al., 2024), InternLM-Math (Ying et al., 2024b), , DeepSeek-Math-Instruct / RL (Shao et al., 2024), Xwin-Math (Li et al., 2024), ToRA (Gou et al., 2024), and NuminaMath (LI et al., 2024). For expert models on theorem proving, we consider several SOTA models with substantial improvements, including reinforcement learning and Monte Carlo search. We include LLM-Step (Welleck and Saha, 2023), GPT-f (Polu and Sutskever, 2020), Lean-STaR (Lin et al., 2024), Hypertree Proof Search (Lample et al., 2022), DeepSeek-Prover (Xin et al., 2024), and InternLM2.5-StepProver (Wu et al., 2024).
@@
Foundation models and proprietary models: We introduce current advanced open-source foundational models, which are Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Mixtral (Jiang et al., 2024). We also introduce proprietary models, which are GPT-4 (OpenAI, 2023), GPT-4o (Hurst et al., 2024) and o1-mini (OpenAI, 2024), as robust competitors.
5 Results and Analysis
5.1 Main Results
This section presents a comprehensive evaluation of CoR-Math-7B’s capabilities in diverse mathematical reasoning tasks. We systematically assess its performance against SOTA models and expert models on specific domains using a zero-shot evaluation setting. Our objective is to demonstrate how multi-paradigm approach enables CoR-Math-7B to achieve superior general mathematical reasoning abilities and computational efficiency, compared to methods based on a single reasoning paradigm.
5.1.1 Comparisons with General-purpose Mathematical Models
Table 1: A overall comparison of CoR-Math-7B with three types of general mathematical reasoners (proprietary, foundational, and general-purpose mathematical models (GMM)) on three mathematical benchmarks. Results are shown for zero-shot (denoted by $\emptyset$) or few-shot settings by default. For the MiniF2F benchmark, we report the best results from the relevant literature with a specified sample budget. Bolded scores are the best performance among all models except for the proprietary one.
Model Model Types MATH GSM8k miniF2F-test Sample Budget ($N$) miniF2F-test Pass@1 Pass@1 Pass@$N$ o1-mini Proprietary 90.0 94.8 1 13.2 GPT-4 Proprietary 42.5 87.1 128 24.6 GPT-4o Proprietary 76.6 90.5 128 25.0 Llama-3.1-8B Foundation 4.2 6.2 128 25.8 Llama-3.1-8B-Instruct Foundation 47.2 76.6 128 23.4 Mistral-7B Foundation 14.3 40.3 1$\times$32$\times$100 22.1 Mixtral-8x7B Foundation 28.4 74.4 1$\times$32$\times$100 23.4 MUSTARD GMM 13.8 27.9 1 7.8 Llemma-7B GMM 18.6 41.0 1$\times$32$\times$100 26.2 DeepSeek-Math-7B-Base GMM 11.8 22.2 1$\times$32$\times$100 28.3 InternLM2-Math-7B-Base GMM 21.5 49.2 1$\times$32$\times$100 30.3 InternLM2-Math-Plus-7B GMM 53.0 85.8 1$\times$32$\times$100 43.3 CoR-Math-7B GMM 66.7 88.7 128$\times$1 52.9 32$\times$100 59.4 128$\times$128 66.0
@@ 5.1.1 Comparisons with General-purpose Mathematical Models
To explore the versatility of CoR-Math-7B and assess its comprehensive mathematical reasoning capabilities, we compare it with widely-used models and SOTA general mathematical models. As shown in Table 1, CoR-Math-7B achieves the best performance across three challenging benchmarks in a zero-shot setting. These results collectively highlight the strong comprehensive mathematical reasoning abilities of our model. The findings are detailed below: (1) arithmetic computation task: our CoR-Math-7B achieves optimal results across all arithmetic computation benchmarks. For example, it exceeds InterLM2-Math-Plus-7B by a $13.7\%$ absolute improvement on the MATH dataset. (2) theorem proving task: our CoR-Math-7B achieves the best zero-shot result on MiniF2F, even surpassing the proprietary GPT-4o model by a $41\%$ absolute increase in few-shot setting. (3) zero-shot vs few-shot: The zero-shot results of CoR-Math-7B outperform all few-shot results from other models. For instance, the zero-shot performance of CoR-Math-7B demonstrates a $14.0\%$ absolute improvement over the few-shot performance of DeepSeek-math-7B-Base on the MiniF2F dataset. When testing on the MiniF2f dataset revealed limited proficiency in Qwen2.5-Math (Yang et al., 2024) in Lean 4, indicating a focus on arithmetic computation over theorem proving. Therefore, we classify Qwen2.5-Math as an arithmetic computation model rather than a general mathematical model and compare it in Section 5.1.3. These findings indicate the strong generalization capability and comprehensive mathematical reasoning ability of the CoR framework. This suggests NLR descriptions and SR verification rehearse precise mathematical reasoning for AR. In summary, the results indicate that the CoR framework can handle a diverse range of complex reasoning challenges, and leveraging the synergy between different paradigms can enhance general reasoning capabilities.
5.1.2 Comparisons with Expert Models on Theorem Proving
Table 2 shows CoR-Math-7B’s exceptional zero-shot performance on the MiniF2F, achieving $66.0\%$ accuracy without demonstrations while competing methods require few-shot examples. Its zero-shot performance is particularly noteworthy when considering computational efficiency.
@@ 5.1.2 Comparisons with Expert Models on Theorem Proving
Table 2: The zero-shot ($\emptyset$) performance of CoR-Math-7B and the few-shot performance of optimized models on the MiniF2F benchmark. These models incorporate optimizations related to theorem proving. The highest scores are in bold.
Model Sample Budget $N$ MiniF2F LLMStep 1$\times$32$\times$100 27.9 GPT-f 64$\times$8$\times$512 36.6 Hypertree Proof Search 64$\times$5000 41.0 Lean-STaR 64$\times$1$\times$50 46.3 DeepSeek-Prover-V1.5-Base 6400 42.2 DeepSeek-Prover-V1.5-SFT + RMaxTS 4$\times$6400 56.3 DeepSeek-Prover-V1.5-SFT + RMaxTS 32$\times$6400 60.2 DeepSeek-Prover-V1.5-RL + RMaxTS 4$\times$6400 59.6 DeepSeek-Prover-V1.5-RL + RMaxTS 32$\times$6400 63.5 InternLM2.5-StepProver 64$\times$32$\times$100 54.5 InternLM2.5-StepProver-BF 1$\times$32$\times$600 47.3 InternLM2.5-StepProver-BF 256$\times$32$\times$600 59.4 InternLM2.5-StepProver-BF+CG 2$\times$32$\times$600 50.7 InternLM2.5-StepProver-BF+CG 256$\times$32$\times$600 65.9 CoR-Math-7B 128$\times$128 66.0
@@
For instance, under similar computational constraints, our CoR-Math-7B surpasses DeepSeek-Prover-V1.5-RL + RMaxTS by $6.4\%$ absolute points ($66.0\%$ vs $59.6\%$ with $128\times 128$ vs $4\times 6400$) and InternLM2.5-StepProver-BF+CG by $15.3\%$ absolute points ($66.0\%$ vs $50.7\%$ with $128\times 128$ vs $2\times 32\times 600$). Compared to significantly larger sample sizes, CoR-Math-7B retains its competitive edge. For example, zero-shot CoR-Math-7B outperforms few-shot InternLM2.5-StepProver-BF+CG by $0.1\%$ in absolute score, despite considering approximately $300$ times more potential solutions. These results highlight the effectiveness of multi-paradigm reasoning and demonstrate how CoR efficiently searches paradigm-based solution spaces.
5.1.3 Comparisons with Expert Models on Arithmetic Computation
Table 3 shows the significant competitiveness of our CoR-Math-7B when compared to expert models on arithmetic computation. Compared to methods like ToRA and NuminaMath that use code as a tool without full reasoning, our CoR-Math-7B with multi-paradigm reasoning outperforms them on all benchmarks. For instance, on the competition-level MATH dataset, CoR-Math-7B achieves an $11.4\%$ absolute improvement over the tool-integrated NuminaMath-7B-TIR. This reinforces the effectiveness of complete code-based reasoning compared to the fragmented interleaving of natural language and Python code.
To further explore the relationship between resource utilization (training sample size) and performance, we fit a quadratic function to the current SOTA models and conduct a Pareto frontier analysis (Branke et al., 2008). Based on Table 3, we generate the fitting Pareto curve shown in Figure 1 (b), which is based on all reasoning methods. Our CoR-Math-7B achieves superior performance compared to the optimal performance of single-paradigm reasoning methods with similar data volume. This suggests the existence of a new optimal curve based on multi-paradigm reasoning methods, which outperforms the curve of single-paradigm approaches.
@@ 5.1.3 Comparisons with Expert Models on Arithmetic Computation
5.2 Ablation Study
We conduct a series of ablation studies, providing insights into the key factors that drive the performance of our multi-paradigm reasoning framework.
5.2.1 Impact of Stages in PPT Method

### v4

@@ 3.3 Training
Inspired by recent advances (LI et al., 2024), we introduce the Progressive Paradigm Training (PPT) strategy, enabling LLMs to gradually master diverse reasoning paradigms. As shown in Fig.˜3 (b), this framework expands the model’s reasoning abilities by sequentially introducing different paradigms at each stage. Each training stage uses a different combination of reasoning paradigms. In stage ➀, given the dominance of NL in the pre-training data of language models (Dubey et al., 2024), we create Numina-CoT as an initialized teaching stage. This is done by modifying the original Numina-CoT dataset (LI et al., 2024) according to our universal text template, enabling the model to learn to use NL to solve complex mathematical problems. Based on the question $x$, the model performs reasoning $\tau_{\text{NLR}}$ and generates the answer $y$. The generated sequence is $z=[x]\tau_{\text{NLR}}y$, where $[x]$ represents the inputs after tokenization. For simplicity, we first consider the loss function for a single sample:
@@
where $\theta$ represents the model parameters, $z_{t}$ is the $t$-th token in the generated sequence, and $z_{<t}$ indicates all tokens before the $t$-th token in the generated sequence. In stage ➁, considering a certain proportion of code corpora in the pre-training data, we expand the training dataset to include two paradigms: NLR and AR. Similar to Numina-CoT, we modify the original Numina-TIR to create Numina-TIR. With this modified dataset, the generated sequence is $z=[x]\tau_{\text{NLR}}\tau_{\text{AR}}y$. After this stage, the model can handle problems that require precise answers. In stage ➂, we further expand the training data to three reasoning paradigms by utilizing the MPM dataset, where $z$ is $x]\tau_{\text{NLR}}\tau_{\text{AR}}\tau_{\text{SR}}y$. After full stages, the trained CoR-Math-7B model not only masters NLR and AR but also can perform rigorous logical SR.
Unlike traditional curriculum and incremental learning, which sequencing tasks by increasing difficulty within a single paradigm or accumulating knowledge, PPT progressively integrates fundamentally distinct reasoning paradigms, transitioning from familiar to unfamiliar, while fostering synergy across paradigms.
@@ 3.4 Inference
We present an inference method combining variable reasoning depth and multi-paradigm sampling to solve comprehensive mathematical tasks.
Prompts with variable reasoning depth. In the zero-shot inference phase, CoR-Math-7B exhibits proficiency in multi-paradigm reasoning, enabling adjustable reasoning depths by modifying prompts based on task requirements. Initially, the model is prompted to conduct NLR, thereby activating a wide range of knowledge patterns associated with natural language from the pre-training corpus. Subsequently, the inference method is tailored to the specific problem type. For example, in theorem proving, the model switches to SR, which is more suitable for formal deduction. Since the SR output is structured in Lean 4, the relevant proof segment can be extracted as the final solution. In arithmetic computations, the model first employs NLR, followed by SR for logical coherence, and concludes with AR for precise calculations. A summary box is utilized to present the final results. The paradigm-switching behavior depends on the presence of prompts and can be categorized into instruction-followed reasoning and instruction-free reasoning. This section primarily presents results from the former. Instruction-free reasoning, which explores how the model autonomously switches paradigms without prompt guidance, will be demonstrated in  Sec.˜D.1. This adaptable paradigm effectively captures the distinct reasoning patterns inherent in different paradigms and demonstrates flexibility in accommodating a variety of scenarios.
Sequential Multi-Paradigm Sampling (SMPS). Instead of token-level sampling based on tree structures (Qiu et al., 2024; Xiong et al., 2024) within single-paradigm reasoning paths, we purpose sequential paradigm-level sampling method, named SMPS. This allows the model to generate outputs by sampling across different paradigms. For instance, in a two-paradigm reasoning scenario, the model first instantiates $J$ distinct paths for the initial reasoning paradigm $\tau_{1}$.
@@
Subsequently, for each of these $\tau_{1}$ paths, the model instantiates $K$ possible paths for the secondary reasoning paradigm $\tau_{2}$.
@@
Overall, the SMPS method utilizes a combinatorial expansion of reasoning paths to explore a diverse paradigm-based solution space, enhancing the robustness and depth of the reasoning results.
@@ 4.1 Evaluation Setup
Datasets. We conduct extensive experiments to evaluate the model’s mathematical reasoning abilities, including arithmetic computation and theorem proving. The arithmetic computation ability is evaluated on the GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), AMC2023 (AI-MO, 2024b) and AIME2024 (AI-MO, 2024a) datasets. Theorem proving ability is evaluated on the miniF2F test set (Zheng et al., 2022), which features mathematical problems of Olympiad difficulty. Further details are in Sec.˜B.4.
[…]
Table 1: A overall comparison of CoR-Math-7B with five types of general mathematical reasoners (Proprietary, Foundational, and General-purpose Mathematical Models (GMM)) on three mathematical benchmarks. Results are shown for zero-shot (denoted by $\emptyset$) or few-shot settings by default. For the miniF2F benchmark, we report the best results from the relevant literature with a specified sample budget. Bolded scores are the best performance among all models except for the proprietary one.
Metrics. Accuracy is the primary evaluation metric. For arithmetic tasks, we adopt widely-used CoT settings (Wei et al., 2022), rounding answers to the nearest integer. The SymPy library [^1] https://www.sympy.org/ is used for parsing and evaluation. To handle numerical representation variations, the model explicitly states the final answer (LI et al., 2024). For theorem proving, we follow the recent advances (Xin et al., 2024), adapting the miniF2F benchmark from Lean 3 to Lean 4. The pass@$N$ metric evaluates proof correctness within $N$ attempts. Our CoR-Math-7B uses SMPS with NLR and SR, where $N=N_{\text{NLR}}\times N_{\text{SR}}$. Model settings are detailed in Sec.˜B.5.
@@ 4.2 Implementation Details
We fine-tuned widely-used DeepSeekMath-Base 7B (Shao et al., 2024) and Llama-3.1 8B (Dubey et al., 2024) models, employing our PPT method on MPM dataset. Unless otherwise specified, CoR-Math-7B model is based on DeepSeekMath-Base 7B. The details are available in Sec.˜B.3.
@@ 4.3 Baselines
We examine three categories of baseline models, with results reported using CoT prompting.
@@
General-purpose mathematical models. To evaluate CoR’s generalization, we includes Mustard (Huang et al., 2024), DeepSeekMath (Shao et al., 2024), InternLM-Math (Ying et al., 2024b), Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Llemma (Azerbayev et al., 2024).
@@
Task-specific mathematical models. We consider several expert models on mathematical optimization for specific tasks, such as large-scale mathematical data and inference search improvements. The arithmetic experts encompass Qwen2.5-Math (Yang et al., 2024), WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2024), DART-Math (Tong et al., 2024), InternLM-Math (Ying et al., 2024b), , DeepSeekMath-Instruct / RL (Shao et al., 2024), Xwin-Math (Li et al., 2024), ToRA (Gou et al., 2024), and NuminaMath (LI et al., 2024). The theorem proving experts include LLM-Step (Welleck and Saha, 2023), GPT-f (Polu and Sutskever, 2020), Lean-STaR (Lin et al., 2024a), Hypertree Proof Search (Lample et al., 2022), DeepSeek-Prover (Xin et al., 2024), and InternLM2.5-StepProver (Wu et al., 2024).
@@
Foundation and proprietary models. We present open-source foundation models Llama-3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), and Mixtral (Jiang et al., 2024), along with proprietary models o1-mini (OpenAI, 2024), GPT-4 (OpenAI, 2023), and GPT-4o (Hurst et al., 2024).
5 Main Results
5.1 Comparisons with General-purpose Mathematical Models
@@
To evaluate the comprehensive mathematical reasoning abilities of CoR-Math-7B, we compare it with widely-used models and SOTA mathematical models, which are based on single reasoning paradigms. As shown in  Table˜1, CoR-Math-7B outperforms others across five challenging benchmarks in a zero-shot setting, highlighting its strong cross-task versatility. The main findings are: (1) arithmetic computation: CoR-Math-7B achieves the best results, outperforming InterLM2-Math-Plus-7B by an absolute $13.7\%$ on MATH and $2.9\%$ on GSM8K. Compared to Llama-3.1-8B-Instruct, it achieves an increase in correct answers on AMC2023 ($34$ vs $16$) and AIME2024 ($12$ vs $5$). (2) theorem proving: CoR-Math-7B sets a new zero-shot benchmark on miniF2F, even surpassing GPT-4o model by a $41.0\%$ absolute increase in few-shot setting. (3) zero-shot vs few-shot: CoR-Math-7B’s zero-shot performance exceeds all few-shot results from other models, with an absolute $37.7\%$ improvement over DeepSeekMath-7B-Base on miniF2F. These findings indicate the generalization capability and comprehensive mathematical reasoning ability of the CoR framework. This suggests NLR descriptions and SR verification rehearse precise mathematical reasoning for AR. Additionally, we provide qualitative analysis of error cases in  Sec.˜D.2. In summary, CoR effectively handles diverse reasoning challenges, with synergies across paradigms boosting its overall performance.
[…]
@@
Table 2: The zero-shot ($\emptyset$) performance of CoR-Math-7B and the few-shot performance of theorem proving optimized models on the miniF2F benchmark. The highest scores are in bold.
5.2 Comparisons with Theorem Proving Experts
@@
Table˜2 shows CoR-Math-7B’s impressive zero-shot performance on miniF2F, achieving $66.0\%$ accuracy without demonstrations, while competitors require few-shot examples. Its zero-shot performance is noteworthy when considering computational efficiency. For instance, under similar computational constraints, CoR-Math-7B surpasses DeepSeek-Prover-V1.5-RL + RMaxTS by $6.4\%$ absolute points ($66.0\%$ vs $59.6\%$ with $128\times 128$ vs $4\times 6400$) and InternLM2.5-StepProver-BF+CG by $15.3\%$ absolute points ($66.0\%$ vs $50.7\%$ with $128\times 128$ vs $2\times 32\times 600$). Even with larger sample sizes, CoR-Math-7B remains superior, surpassing few-shot InternLM2.5-StepProver-BF+CG by $0.1\%$, despite considering $300$ times more solutions. These results highlight the efficiency of multi-paradigm reasoning and CoR’s ability to explore paradigm-based solution spaces.
@@
5.3 Comparisons with Arithmetic Experts
Since several mathematical reasoners are primarily trained on arithmetic computation rather than formal theorem proving, and often struggle with the latter, we classify them as experts in arithmetic computation to enable a fair comparison. Table˜3 highlights CoR-Math-7B’s strong performance in arithmetic computation, surpassing expert models and achieving a $15\%$ improvement over RL-based methods, with DeepSeekMath-RL-7B as the representative, on the MATH benchmark for arithmetic tasks. Unlike methods such as ToRA and NuminaMath, which rely on code without complete reasoning, CoR-Math-7B leverages multi-paradigm reasoning to outperform them across all benchmarks. For example, on the MATH dataset, CoR-Math-7B achieves an $11.4\%$ absolute improvement over the tool-integrated NuminaMath-7B-TIR, underscoring the effectiveness of complete code-based reasoning over fragmented interleaving of natural language and code. To further explore the relationship between resource utilization (training sample size) and performance, we fit a quadratic function to the current SOTA models and perform a Pareto frontier analysis (Branke et al., 2008), as shown in Fig.˜1 (Details in Sec.˜B.6). CoR-Math-7B outperforms the optimal performance of single-paradigm methods with similar data volume, suggesting a new optimal curve for multi-paradigm reasoning that surpasses single-paradigm approaches.
6 Ablation Study